# XSNAP: An X-ray Supernova Analysis Pipeline with Application to the Type II Supernova 2024ggi

### v1

@@ 3.2 CSM Density Profile Analysis
where $n_{e}$ and $n_{I}$ are in $\rm cm^{-3}$ and $d$ is the distance to the source in $\rm cm^{-1}$. We follow the procedure of D. Brethauer et al. (2022) to calculate the unshocked CSM density profile, which goes as:
@@
where $\mu_{e}$ and $\mu_{I}$ are the mean molecular weight of electron and ion respectively; $m_{p}$ is the proton mass; $\mathrm{V_{FS}}=\frac{4\pi}{3}f\left(R_{\rm out}^{3}-R_{\rm in}^{3}\right)$ is the emitting forward shock (FS) volume, with $R_{\rm in}$ and $R_{\rm out}$ are the inner and outer radius of the shell respectively and $f$ is the filling factor. In this work, we adopt $R_{\rm out}=1.2R_{\rm in}$, $f=1$, $\mu_{e}=1.14$, and $\mu_{I}=1.24$ (assuming solar values). We also estimated the shock radius by $R_{\rm shock}=v_{\rm shock}\times t$, where $t$ is time since explosion and assumed an average FS velocity $v_{\rm shock}=10^{4}\ \rm km\ s^{-1}$. The mass-loss rate was then estimated by fitting the density profile to $\rho_{CSM}(r)=\frac{\dot{M}}{4\pi r^{2}v_{\rm wind}}$ with the emcee package. Incorporating $v_{\rm wind}=20\ \rm km\ s^{-1}$, we inferred a mass-loss rate of $\dot{M}=(6.2\pm 0.2)\times 10^{-5}\ \rm M_{\odot}\ yr^{-1}$ from the detected emission for the last 117 years before explosion, which lies between the upper-limit of the mass-loss rates of Type IIP SNe $\sim 10^{-5}\ \rm M_{\odot}\ yr^{-1}$ and the lower-limit of the mass-loss rates of Type IIn SNe $\sim 10^{-3}\ \rm M_{\odot}\ yr^{-1}$ (V. V. Dwarkadas, 2025). The density profile and its comparison to SN 2023ixf and density profile models by W. V. Jacobson-Galán et al. (2024a) are shown in Figure 3.
4 The Pipeline
Each X-ray observatory, such as CXO, XMM-Newton, and Swift-XRT, and Nuclear Spectroscopic Telescope Array (NuSTAR), employs mission-specific data formats, calibration pipelines, and analysis software. Consequently, this hinders reproducibility in analyzing X-ray observations as astronomers need to work out distinct guidelines for each mission. While there have been past efforts to streamline the process, such as YAXX [^1] https://cxc.harvard.edu/contrib/yaxx (Yet Another X-ray Xtractor; T. Aldcroft 2006), which supported both CXO and XMM-Newton data, these tools are no longer actively maintained. More recent softwares, like XGA [^2] https://xga.readthedocs.io/en/latest/ (X-ray: Generate and Analyse; D. J. Turner et al. 2022) or BXA [^3] https://johannesbuchner.github.io/BXA/ (Bayesian X-ray Analysis; J. Buchner et al. 2014), offer modern functionality but are limited to a single instrument (e.g., XMM-Newton), not designed as end-to-end pipelines from spectral extraction, or not specifically developed for SNe, even though some may still be adaptable for such studies. At present, there is no open-source, Python-based pipeline that unifies spectral extraction and analysis of SN X-ray spectra from CXO, XMM-Newton, Swift-XRT, and NuSTAR. Consequently, we developed XSNAP (X-ray Supernova Analysis Pipeline), an open-source Python package intended to unify end-to-end X-ray data processes from spectral extraction to CSM analysis.
XSNAP is organized into two stages: spectrum extraction and spectral modeling. XSNAP is designed to streamline all the processes utilized in Sections 2 and 3, e.g., source detection, deriving count-rates, spectral fitting and deriving unshocked CSM densities. The spectrum extraction component standardizes the procedures for each observatory described in Section 2, implementing them as dedicated Python Command-line Interface scripts. In addition, a script for generating source and background regions is provided, making use of ximage and (optionally) SAOImage DS9. For NuSTAR spectrum extraction process, we follow the standard guide from NuSTAR Data Analysis Software (NuSTARDAS [^4] https://heasarc.gsfc.nasa.gov/docs/nustar/analysis/nustar_quickstart_guide.pdf) (F. A. Harrison et al., 2013) inside HEASoft. Similarly, we developed the spectrum and CSM analysis sector by wrapping all the processes from Section 3 to each separate Python module. We then packaged the pipeline in a Python Package Index (PyPI) environment and published it to Github [^5] https://github.com/fercananything/XSNAP/ and PyPI [^6] https://xsnap.org/. Accordingly, users can streamline the whole analysis process by (i) making the source and background regions and (ii) extracting and reducing raw spectrum through the Python CLI scripts and (iii) analyzing spectrum from the Python API modules.
5 Discussion
@@
Detected X-ray observations of SN 2024ggi up to $\sim 85$ days after explosion constrain both the shock properties and the progenitor activity before core collapse. SN 2024ggi has a progenitor mass-loss rate of $(6.2\pm 0.2)\times 10^{-5}\ \rm M_{\odot}\ yr^{-1}$ for a steady-state wind velocity of $v_{\rm w}=20\ \rm km\ s^{-1}$ for the last 117 years before first light based on the phases of detected X-ray emission. This mass-loss rate is higher than what is estimated for most SNe IIP with X-ray observations (e.g., see I. Irani et al. 2024; V. V. Dwarkadas 2025), but comparable to the SN 2023ixf progenitor mass-loss rate of $\sim 10^{-4}\ \rm\ M_{\odot}\ yr^{-1}$ (A. J. Nayana et al., 2025; W. V. Jacobson-Galán et al., 2025a; K. A. Bostroem et al., 2025) at shock radii $10^{15}$ cm. Based on Figure 6 in J. Fuller & D. Tsuna (2024), $6.2\times 10^{-5}\ \rm M_{\odot}\ yr^{-1}$ corresponds to a RSG progenitor mass of $\sim 20\ \rm M_{\odot}$. This estimate is higher than masses typically inferred from pre-explosion imaging and nebular-phase spectroscopy for CSM-interacting SNe II (e.g., C. D. Kilpatrick et al. 2025; W. V. Jacobson-Galán et al. 2025b), including SN 2024ggi (L. Ferrari et al., 2025; D. Xiang et al., 2024). Additionally, the X-ray luminosity peaks around the first $\sim 5$ days after first light and decays as $L_{X}\propto t^{-0.99}$, which is approximately aligned with a steady-state wind mass-loss who has a luminosity decay of $L\propto t^{-1}$.
@@
A temporal evolution of the intrinsic neutral hydrogen column $(\rm NH_{\rm int})$ provides evidence for a dense, confined CSM (see Table 1). Although our results agree with W. V. Jacobson-Galán et al. (2024a) that a dense, confined CSM is present, the densities inferred from the optical spectra are higher than those derived from the X-rays at $5\times 10^{14}$ cm. This discrepancy, also observed in SN 2023ixf, may arise from CSM asymmetry or clumping (A. Singh et al., 2024; A. J. Nayana et al., 2025; S. S. Vasylyev et al., 2025). At early times $(\delta t\lesssim 16\ \rm days)$, we found $\rm NH_{\rm int}>10^{22}\ \rm cm^{-2}$, indicating significant X-ray absorption in a dense medium extending up to $r\lesssim 10^{15}\ \rm cm$, before this material overtaken by the SN shock at $\delta t\gtrsim 11\ \rm days$. All of these results were produced with our pipeline, XSNAP, which streamlines the entire workflow from raw counts to CSM analysis. Beyond this application to SN 2024ggi, XSNAP is broadly applicable: its data reduction and spectral modeling modules can be used for other X-ray sources, while its CSM analysis is specialized for SNe.
6 Conclusion
@@
  • SN 2024ggi shows evidence for a dense CSM $(\rm NH_{\rm int}>10^{22}\ \rm cm^{-2})$ up to $\sim 16$ days after explosion, confined within $r\lesssim 1.4\times 10^{15}\ \rm cm$.
@@
$F.$ acknowledges support from the Caltech Summer Undergraduate Research Fellowship (SURF) program. W.J.-G. is supported by NASA through Hubble Fellowship grant HSTHF2-51558.001-A awarded by the Space Telescope Science Institute, which is operated for NASA by the Association of Universities for Research in Astronomy, Inc., under contract NAS5-26555. This research has made use of data and/or software provided by the High Energy Astrophysics Science Archive Research Center (HEASARC), which is a service of the Astrophysics Science Division at NASA/GSFC.
@@ Appendix B X-Ray Observation Logs
Instrument Observation Date (yyyy/mm/dd) Mid Timea (days) Observation ID Exposure Time (ks) PI
Swift-XRT 2024-04-11 - 2024-04-16 3.33 00045607063, 00016601001, 00016601002, 00016601003, 00016601004, 00016601005, 00016601006, 00016601007, 00016601012, 00016601013, 00016601014, 00016601015, 00016601017, 00016601018, 00016601019, 00045607064, 00045607065 20 Swift, Sand
Swift-XRT 2024-04-17 - 2024-04-20 8.36 00016601020, 00016601022, 00016601023, 00016601024, 00016601025, 00016601026, 00016601027 9 Swift, Sand
@@
Swift‐XRT 2024-05-01 – 2024-05-09 24.49 00016601044, 00016601045, 00016601048, 00016601049, 00016601050, 00016601051, 00016601052, 00016601053 7.6 D. Sand
Swift‐XRT 2024-05-12 – 2024-05-18 34.76 00016601055, 00016601059, 00016601062, 00016601063 5.0 D. Sand
Swift‐XRT 2024-05-21 – 2024-05-26 43.2 00016601064, 00016601065, 00016601066, 00016601067, 00016601068 10.1 D. Sand
@@
Swift‐XRT 2024-06-02 – 2024-06-08 55.7 00016601069, 00016601070, 00016601071 2.4 D. Sand
Swift‐XRT 2024-06-24 – 2024-07-08 81.51 00016601072, 00016601073, 00016601074 6.1 D. Sand
@@
Swift‐XRT 2024-07-15 – 2024-08-05 106.35 00016601075, 00016601076, 00016601077, 00016601078 4.7 D. Sand
Swift‐XRT 2024-08-08 – 2024-08-19 125.21 00016601079, 00016601080, 00016601081 4.9 D. Sand
@@
Swift‐XRT 2024-12-21 – 2025-01-01 260.33 00045607067, 00045607068, 00045607069, 00045607070 5.2 Swift
Swift‐XRT 2025-01-08 – 2025-01-15 276.65 00045607071, 00045607072, 00045607075 4.0 Swift
Swift‐XRT 2025-02-08 – 2025-03-05 316.43 00045607076, 00045607078, 00045607080 3.7 Swift
Swift‐XRT 2025-03-10 – 2025-03-20 338.76 00045607081, 00045607082 2.9 Swift
Notes: a With respect to time of explosion

### v2

@@ IV.2 CSM Density Profile Analysis
where $n_{e}$ and $n_{I}$ are in $\rm cm^{-3}$ and $d$ is the distance to the source in $\rm\mathbf{c}m$. We follow the procedure of D. Brethauer et al. (2022) to calculate the unshocked CSM density profile, which goes as:
@@
where $\mu_{e}$ and $\mu_{I}$ are the mean molecular weight of electron and ion respectively; $m_{p}$ is the proton mass; $\mathrm{V_{FS}}=\frac{4\pi}{3}f\left(R_{\rm out}^{3}-R_{\rm in}^{3}\right)$ is the emitting forward shock (FS) volume, with $R_{\rm in}$ and $R_{\rm out}$ are the inner and outer radius of the shell respectively and $f$ is the filling factor. In this work, we adopted $R_{\rm out}=1.2R_{\rm in}$, $f=1$, $\mu_{e}=1.14$, and $\mu_{I}=1.24$ (assuming solar values). We also estimated the shock radius by $R_{\rm shock}=v_{\rm shock}\times t$, where $t$ is time since explosion and assumed an average FS velocity $v_{\rm shock}=10^{4}\ \rm km\ s^{-1}$. The mass-loss rate was then estimated by fitting the density profile to $\rho_{CSM}(r)=\frac{\dot{M}}{4\pi r^{2}v_{\rm wind}}$ with the emcee package. Incorporating $v_{\rm wind}=20\ \rm km\ s^{-1}$, we inferred a mass-loss rate of $\dot{M}=(6.2\pm 0.2)\times 10^{-5}\ \rm M_{\odot}\ yr^{-1}$ from the detected emission for the last $\sim 117$ years before explosion, which lies between the upper-limit of the mass-loss rates of Type IIP SNe $\sim 10^{-5}\ \rm M_{\odot}\ yr^{-1}$ and the lower-limit of the mass-loss rates of Type IIn SNe $\sim 10^{-3}\ \rm M_{\odot}\ yr^{-1}$ (V. V. Dwarkadas, 2025). This lookback time can be treated as an approximation as they heavily rely on the wind and shock velocities, which both are assumed. The density profile and its comparison to SN 2023ixf and density profile models by W. V. Jacobson-Galán et al. (2024a) are shown in Figure 3.
V Discussion
@@
Detected X-ray observations of SN 2024ggi up to $\sim 85$ days after explosion constrain both the shock properties and the progenitor activity before core collapse. SN 2024ggi has a progenitor mass-loss rate of $(6.2\pm 0.2)\times 10^{-5}\ \rm M_{\odot}\ yr^{-1}$ for a steady-state wind velocity of $v_{\rm w}=20\ \rm km\ s^{-1}$ for the last $\sim 117$ years before first light based on the phases of detected X-ray emission. This mass-loss rate is higher than what is estimated for most SNe IIP with X-ray observations (e.g., see I. Irani et al. 2024; V. V. Dwarkadas 2025), but comparable to the SN 2023ixf progenitor mass-loss rate of $\sim 10^{-4}\ \rm\ M_{\odot}\ yr^{-1}$ (A. J. Nayana et al., 2025; W. V. Jacobson-Galán et al., 2025a; K. A. Bostroem et al., 2025) at shock radii $10^{15}$ cm. Based on Figure 6 in J. Fuller & D. Tsuna (2024), $6.2\times 10^{-5}\ \rm M_{\odot}\ yr^{-1}$ corresponds to a RSG progenitor mass of $\sim 20\ \rm M_{\odot}$. This estimate is higher than masses typically inferred from pre-explosion imaging and nebular-phase spectroscopy for CSM-interacting SNe II (e.g., C. D. Kilpatrick et al. 2025; W. V. Jacobson-Galán et al. 2025b), including SN 2024ggi (L. Ferrari et al., 2025; D. Xiang et al., 2024). Additionally, the X-ray luminosity peaks around the first $\sim 5$ days after first light and decays as $L_{X}\propto t^{-0.99}$, which is approximately aligned with a steady-state wind mass-loss who has a luminosity decay of $L\propto t^{-1}$.
@@
A temporal evolution of the intrinsic neutral hydrogen column $(\rm N_{\rm H,\ int})$ provides evidence for a dense, confined CSM (see Table 1). Although our results agree with W. V. Jacobson-Galán et al. (2024a) that a dense, confined CSM is present, the densities inferred from the optical spectra are higher than those derived from the X-rays at $5\times 10^{14}$ cm. This discrepancy, also observed in SN 2023ixf, may arise from CSM asymmetry or clumping (A. Singh et al., 2024; A. J. Nayana et al., 2025; S. S. Vasylyev et al., 2025). At early times $(\delta t\lesssim 16\ \rm days)$, we found $\rm N_{\rm H,\ int}>10^{22}\ \rm cm^{-2}$, indicating significant X-ray absorption in a dense medium extending up to $r\lesssim 10^{15}\ \rm cm$, before this material overtaken by the SN shock at $\delta t\gtrsim 11\ \rm days$. All of these results were produced with our pipeline, XSNAP, which streamlines the entire workflow from raw counts to CSM analysis. Beyond this application to SN 2024ggi, XSNAP is broadly applicable: its data reduction and spectral modeling modules can be used for other X-ray sources, while its CSM analysis is specialized for SNe.
VI Conclusion
@@
  • SN 2024ggi shows evidence for a dense CSM $(\rm N_{\rm H,\ int}>10^{22}\ \rm cm^{-2})$ up to $\sim 16$ days after explosion, confined within $r\lesssim 1.4\times 10^{15}\ \rm cm$.
@@
F. acknowledges support from the Caltech Summer Undergraduate Research Fellowship (SURF) program. W.J.-G. is supported by NASA through Hubble Fellowship grant HSTHF2-51558.001-A awarded by the Space Telescope Science Institute, which is operated for NASA by the Association of Universities for Research in Astronomy, Inc., under contract NAS5-26555. This research has made use of data and/or software provided by the High Energy Astrophysics Science Archive Research Center (HEASARC), which is a service of the Astrophysics Science Division at NASA/GSFC. This paper employs a list of Chandra datasets, obtained by the Chandra X-ray Observatory, contained in the Chandra Data Collection  https://doi.org/10.25574/cdc.531 (catalog DOI: 10.25574/cdc.531).
@@ Appendix B X-Ray Observation Logs
Instrument Observation Mid Observation ID Exposure PI
Date TimeaaWith respect to time of explosion. Time (ks)
(yyyy/mm/$d$) (days)
Swift-XRT 2024-04-11 – 3.33 00045607063, 00016601001, 00016601002, 20 Swift,
2024-04-16 00016601003, 00016601004, 00016601005, D. Sand
00016601006, 00016601007, 00016601012,
00016601013, 00016601014, 00016601015,
00016601017, 00016601018, 00016601019,
00045607064, 00045607065
Swift-XRT 2024-04-17 – 8.36 00016601020, 00016601022, 00016601023, 9 Swift,
2024-04-20 00016601024, 00016601025, 00016601026, D. Sand
00016601027
@@
Swift‐XRT 2024-05-01 – 24.49 00016601044, 00016601045, 00016601048, 7.6 D. Sand
2024-05-09 00016601049, 00016601050, 00016601051,
00016601052, 00016601053
Swift‐XRT 2024-05-12 – 34.76 00016601055, 00016601059, 00016601062, 5.0 D. Sand
2024-05-18 00016601063
Swift‐XRT 2024-05-21 – 43.2 00016601064, 00016601065, 00016601066, 10.1 D. Sand
2024-05-26 00016601067, 00016601068
@@
Swift‐XRT 2024-06-02 – 55.7 00016601069, 00016601070, 00016601071 2.4 D. Sand
2024-06-08
Swift‐XRT 2024-06-24 – 81.51 00016601072, 00016601073, 00016601074 6.1 D. Sand
2024-07-08
@@
Swift‐XRT 2024-07-15 – 106.35 00016601075, 00016601076, 00016601077, 4.7 D. Sand
2024-08-05 00016601078
Swift‐XRT 2024-08-08 – 125.21 00016601079, 00016601080, 00016601081 4.9 D. Sand
2024-08-19
@@
Swift‐XRT 2024-12-21 – 260.33 00045607067, 00045607068, 00045607069, 5.2 Swift
2025-01-01 00045607070
Swift‐XRT 2025-01-08 – 276.65 00045607071, 00045607072, 00045607075 4.0 Swift
2025-01-15
Swift‐XRT 2025-02-08 – 316.43 00045607076, 00045607078, 00045607080 3.7 Swift
2025-03-05
Swift‐XRT 2025-03-10 – 338.76 00045607081, 00045607082 2.9 Swift
2025-03-20